OpenStudy (anonymous):

Identify the given surface x^2+4y^2-10z=0

OpenStudy (anonymous):

level curves are x=0 : z= (2/5)*y^2 (parabola in the yz) y=0 : z= (1/10) *x^2 (parabola in the xz plane) z=C x^2 +4y^2 =10C (ellipses parallel to the xy plane) so it's a...?